OpenStudy (anonymous):

Find the missing measure for a right circular cone given the following information. Find h if r = 5 and V = 100. 4 12 30

OpenStudy (michele_laino):

please what is the formula expressing the volume V of a right circular cone?

OpenStudy (michele_laino):

do you know it?

OpenStudy (michele_laino):

@Luka172

OpenStudy (anonymous):

i think its V =1/3 pi r^2 h

OpenStudy (michele_laino):

perfect!

OpenStudy (michele_laino):

Now, multiply both side of your formula by 3, and subsequently divide your resultant formula by pi*r^2, write your result please!

OpenStudy (michele_laino):

please, first multiply both sides of the formula above that you have wrote by 3, the write your answer, please!

OpenStudy (anonymous):

300=25pi/3 i dont know

OpenStudy (michele_laino):

I write the first step: \[3V=\pi*r ^{2}*h\]

OpenStudy (michele_laino):

now, please divide my formula above by (pi*r^2), please do not substitute your numerical data, we will do it later!

OpenStudy (michele_laino):

please you have to divide both sides of my ormula by (pir^2)

OpenStudy (michele_laino):

right side, will become: \[\frac{ \pi*r ^{2}*h }{ \pi*r ^{2} }=h\] and left side?...please write your answer

OpenStudy (anonymous):

im stupid i dont know can u put it in slow turms

OpenStudy (michele_laino):

your right side is: \[\pi*r ^{2}*h\] now when you divide it by (pi*r^2), it becomes: \[\frac{ \pi*r^{2}* h}{ \pi*r ^{2} }=h\]

OpenStudy (michele_laino):

left side will become: \[\frac{ 3V }{ \pi*r ^{2} }\]

OpenStudy (anonymous):

i got that but how do i put the problem in the calculator

OpenStudy (michele_laino):

so finally yuo got: \[h=\frac{ 3V }{ \pi*r ^{2} }\]

OpenStudy (michele_laino):

please insert our numerical data at right side of the above formula, and you wil get your answer Please you cn use this approximation: \[\pi \approx 3\]

OpenStudy (michele_laino):

so:\[h=\frac{ 3*100 }{ 3*25 }=4\]

OpenStudy (michele_laino):

so the first option is the ight answer!

OpenStudy (anonymous):

but what about 12

OpenStudy (michele_laino):

why 12?